THE VISIBLE SPECTRUM OF Si_{3}

XIUJUAN ZHUANG, \underline{TIMOTHY C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; N. REILLY, D. KOKKIN and M. C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, Cambridge, Massachusetts 02138, USA; J. F. STAN-TON , Chemistry Department and Biochemistry, U. of Texas, Austin, TX 78712, USA; T. D. CRAW-FORD and B. FORTENBERRY, Chemistry Department, Virgina Tech, Blackbury VA 24061, USA; J. P. MAIER , Department of Chemistry, University of Basel, Basel, Switzerland.

We report the first recording of the visible spectrum of gas-phase Si_{3}. A cold molecular beam sample was produced by skimming the output of a pulsed discharge source. The 545-490 nm spectral region was examined using both massselected REMPI and pulsed dye laser excitation with LIF detection. Dispersed fluorescence and lifetime measurements of numerous bands were recorded. The spectrum has an origin at 18600 cm^{−1} and a progression in the symmetric stretch with a harmonic frequency of 445 cm^{−1}. The bands are assigned to the 1^{3}A_{1}′′- ãA_{2}′ transition of the D_{3h} isomer based upon new and previous^{a} predictions. A vibrational progression observed in the dispersed fluorescence having a spacing of 505 cm^{−1} is in agreement with previous ZEKE studies^{b}. An additional vibrational progression observed in dispersed fluorescence having a spacing of 173 cm^{−1} is in agreement with the bending frequency for the X̃^{1}A_{1} state of the C_{2v} isomer estimated from the pure rotational spectrum.^{c}

aP. Garcia-Fernandez, J. Boggs and J. P. Stanton J. Chem. Phys. 126, 074305 (2007)

bC. C.Arnold and D. M. Neumark J. Chem. Phys. 100, 1797 (1994)

cM. C.McCarthy and P. Thaddeus Phys.Rev. Lett. 90, 213003 (2003)